# TRAIL: Trace Reasoning and Agentic Issue Localization

### v1

@@ 4.3 Dataset Analysis
Figure 2: Distribution of Errors per Trace
As a result of the post annotation review, we observed that a total of 144 traces were found to contain errors and 114 traces in GAIA and 30 from the SWE Bench contained at least one error. Our annotators identified a total of 841 unique errors in all traces with an average of 5.68 and a median of 5 errors per trace (see Figure 2).
Further breakdown in Figure 3 shows that errors span a wide range of categories, with most errors belonging to the Output Generation category. Formatting Errors or Instruction Non-compliance errors in particular account for 353 of the 841 errors (or nearly 42%). Meanwhile, there do not appear to be many instances of System Execution Errors. Although this categorical imbalance stems directly from the type of traces we use as input, we believe it highlights two key considerations when evaluating modern agentic pipelines. Firstly, the high concentration of Output Generation errors indicates that despite best efforts at prompt-engineering, LLM systems have a hard time with higher-level reasoning and understanding the parameters of the task they are given. Secondly, the large number of categories with very few errors are several times, catastrophic for the progress of the system. For example, recoverability is severely hindered for cases where an API fails as opposed to when systems deviate from goals or misinterpret tool outputs. Despite appearing infrequently, these error categories are the most important to identify and evaluate for practical agentic applications.
(a) Category Annotations
(b) Parent Category Annotations
(c) Root Category Annotations
Figure 3: Error Category Distributions in TRAIL as per different levels of the taxonomy
Most of the errors in our dataset fall in either the high or medium categories (6(a)). Model hallucinations and resource management issues have a high impact on the agent behavior, whereas nearly 44% of the Output Generation errors are classified as Low impact (6(b)). This further emphasizes the importance of a classification schema which is able to capture "high-value" infrequent or rare behaviors. We view the ability to identify and classify each of the samples in the long-tail of high-errors into a specific category as a key feature of our taxonomy.
4.4 Evaluation Setup
@@ 5 Results
In the following subsections, we analyze the research questions below:
  1. 1.
    How is TRAIL performance affected by long context ability of the LLM? How many of the input instances exhaust the LLM’s context window? How does it vary with length of the trace? We answer these in §5.1.1 §5.1.2, and §5.1.3.
  2. 2.
    Does TRAIL benefit from more reasoning? We answer this question in §5.1.4 and §5.1.5.
  3. 3.
    Which error categories are easier to predict for models when predicting TRAIL over traces? On which particular ones is performance substantively worse for non-reasoning models? We answer these questions in §5.1.6
@@ 5.1.1 Task Difficulty - Context Length and Generation Horizon
As we can see in Table 2, the distribution of input token lengths that need to be ingested in the form of raw trace jsons to perform our task cuts close to the input context limit of several LLMs - with the maximum input trace length always being twice longer than the input length limit, and even the mean itself sometimes going over. Furthermore, even the typical output token length horizon the LLMs need to generate for the task exceeds the 1K token mark on average, with the maximum being $\approx$3.7K at the least. Besides being a significant % of the maximum output token length, this indicates the challengingly long generation horizon generating for TRAIL entails.
@@
Table 3: Correlations between Input Length & Performance
5.1.2 Long Context Ability and its effect on Model Performance
We compare how the models in Table 1 rank based on their aggregate performance on TRAIL vis-a-vis the relative ranking of the subsets of these models that occur on the freshly [per May 2025] updated long-context benchmark leaderboards Longbenchv2 and fiction.live’s LongContextBench Bai et al. (2024b); Ficlive (2025), and notice this differs for only one model (o3 being third best rather than best on the latter). We defer the complete detail of the rankings read off these leaderboards to §A.1.1.
5.1.3 Performance as a Function of Input Length
We find all performance metrics to be anti-correlated with input token length, as detailed in Table 3. This supports the hypothesis that longer input raw traces increase the difficulty of TRAIL for models.
@@ 5.1.4 Reasoning vs Non-Reasoning Models
From Table 1, we see all post-2024 reasoning models (i.e. save o1) outperforming non-reasoning ones on both Error category F1 and Location Accuracy. On Joint Accuracy i.e. getting both error category and location right, the gap between the two families is larger. Reasoning models other than o1 perform at 1.5-8 times the best performing non-reasoning model in terms of joint accuracy.
@@ 5.1.5 Does Reasoning Effort Matter?
To ablate for the effect of extent of reasoning in a controlled fashion, we run experiments for the same model (o3) with different levels of effort - "high", "medium" and "low", as supported by OpenAI’s API through its reasoning.effort parameter. We observe a steady decrease in all three metrics with decrease in the extent of reasoning. For instance, Category F1 goes from 0.296 for "high" $\rightarrow$ 0.277 for "medium" $\rightarrow$ 0.264 for "low".
These observations empirically validate the hypothesis that TRAIL benefits from the model being able to expend reasoning effort at test time, and the superior numbers reasoning models we observe on TRAIL (noted in §5.1.4) are not merely courtesy better post-training or pre-training. We defer the complete reasoning effort ablation results to Appendix §A.2.
@@ 5.1.6 Performance Across Categories
Figure 5: Heatmap visualizing Error Category F1 across models; categories are ordered left to right based on their support
  1. 1.
@@
    • Context Handling Failures is one of the most challenging categories, with most models showing an F1 score of 0.00, suggesting this type of error requires sophisticated reasoning. The only model that performs passably atleast is claude-3.7-sonnet at 0.18
    • Tool Selection Errors has poor performance across most models (0.00-0.08 range), except for gemini-2.5-pro (0.26), claude-3.7-sonnet (0.27) and o3 (0.53), indicating this complex error type is difficult to detect.
    • Task Orchestration has poor performance across most models (0.00-0.08 range) apart from gemini-2.5-flash which does significantly better on it with a F1 of 0.47
  2. 2.
@@
    • Goal Deviation shows vivid performance differences, with gemini-2.5-pro/flash doing relatively best (0.70/0.41), claude-3.7-sonnet and o3 performing moderately (0.31,0.24); and lastly o1 and the other non-reasoning models performing the lowest ($\leq 0.05$)
    • For Poor Information Retrieval, the two gemini models gemini-2.5-pro/flash are distinctively better (0.50/0.53) than the others which have F1 below 0.30. This points to an advantage these models might have at diagnosing incorrect conditioning on the context as a whole.
  3. 3.
@@
    • Language-Only (a type of Hallucination) errors are relatively well-detected across all models (0.14-0.59), suggesting that predicting this error category is more accessible even without reasoning.
    • Formatting Errors shows some interesting non-monotonicity and performance bands w.r.t - gpt-4.1 (0.43) and gemini-2.5 models (0.44-0.57) perform well, while o1, o3 and claude-3.7-sonnet occupy a lower band (0.23-0.31). Surprisingly, both gpt-4.1 and o1 outdo o3 here inspite of being non-reasoning and an earlier generation respectively.
  4. 4.
    Model-Specific Observations
    • gemini-2.5-pro is clearly the top performer, with particularly strong relative performance on: Goal Deviation (0.70), Poor Information Retrieval (0.50) and Tool Output Misinterpretation (0.67) and Environment Setup Errors (0.57)
    • gpt-4.1’s performance varies highly sensitive to category, doing very well or moderately on some (Instruction Non-compliance, Language-only, Formatting Errors, Resource Abuse), while going below the 0.10 mark and even hitting a complete zero on some others (Goal Deviation, Tool Selection Errors, Task Orchestration, Tool-Related Hallucinations and Context Handling Failures).
@@ 6 Conclusion
In this work, we devised TRAIL, a new taxonomy for agentic error classification. We accompanied this taxonomy with an expert curated dataset of 148 agentic problem instances consisting of 841 unique error derived from two publicly available datasets: GAIA and SWE Bench. We demonstrated the inability of SOTA models to act as LLM Judges, with the best performing model on our dataset (Gemini 2.5-pro) only able to achieve 18% joint accuracy on GAIA, and 5% joint accuracy on SWE Bench. Meanwhile, three of the eight models we test are unable to handle the complexity of TRAIL and fail to solve the task due to context-length limitations. The benchmark performance of the SOTA models on our dataset shows that these models are unable to systematically evaluate complex traces generated by agentic systems. This problem stems from the fundamental complexity of large agentic systems and finite context limitations of LLMs. We require a new framework in order to systematically and scalably evaluate agentic workflows.
Limitations & Future Work
The TRAIL dataset and taxonomy are primarily focused on text-only inputs and outputs but recent advancements in multi-modal agentic systems require careful extension of the taxonomy to handle multimodal errors arising from newer categories such as multimodal tool use. One additional limitation of TRAIL is the large number of tail categories with very few examples. It is important to ensure correctness of LLM-Judges on these categories due to the high-impact nature of the failures. Future research work can look into synthetic data generation for high-impact, low-occurrence categories by systematically modifying existing traces to induce catastrophic irrecoverable failures within the LLM context.
@@
While curating this dataset, we ensure that annotators are only selected based on their age (18+) and their expertise in the computer science field. Annotator selection was not based on nationality, language, gender or any other characteristic apart from these two criteria. We pay annotators a total of \$12.66 per trace where each trace takes 30-40 minutes to annotate. We ensure that the traces do not contain any PII or any explicit or biased content by manually verifying traces before forwarding these to annotators.
@@ A.1 Prompt Structure
A.1.1 Long Context Leaderboard Rankings vs TRAIL
@@
A.2 Reasoning Effort Ablations
@@ A.3 Span Statistics
A.3.1 Annotation Rubric
@@
A.4 Correlation scores for Rubrics
@@
A.5 Distribution of Impact Levels in TRAIL instances
@@
A.6 Agent Orchestrations for TRAIL

### v2

@@ 4.3 Dataset Analysis
Following the post-annotation review, we found errors in 114 GAIA traces and 30 from SWE Bench. As shown in Figure 3, these errors cover various categories, with most falling under Output Generation. Specifically, Formatting Errors and Instruction Non-compliance make up 353 of 841 total errors—nearly 42%. In contrast, System Execution Errors are rare. This categorical imbalance highlights two important considerations for evaluating agentic pipelines. First, the prevalence of Output Generation errors suggests that current LLM systems struggle with high-level reasoning and understanding task parameters, even with careful prompt-engineering. Second, although infrequent, errors in categories like API failures can be catastrophic and are critical to detect, as they are often difficult to recover from, unlike errors due to goal deviation or tool misinterpretation. Most errors in our data are high or medium impact (6(a)). While model hallucinations and resource management issues greatly affect agent behavior, about 44% of Output Generation errors are low impact (6(b)). This underscores need for a classification scheme that includes rare but significant error types. A key feature of our taxonomy is ability to categorize well such long-tail, high-impact errors.
4.4 Summary of Evaluation Setup
For empirically evaluating and comparing LLM performance on TRAIL we choose the following LLMs — gpt-4.1, o1, o3, gemini-2.5 (both pro+flash), claude-3.7-sonnet and LLama-4 (both Maverick+Scout). We defer detailed discussion of more evaluation setup specifics to A.3
@@ 5 Results
In §5.1, we analyze the research questions below:
  • How does long context reasoning affect TRAIL performance? How many inputs exceed the LLM’s context window? How does trace length impact this? We address these in §5.1.1 §5.1.2, and §5.1.3.
  • Does TRAIL benefit from more reasoning? We explore this in §5.1.4 and §5.1.5.
  • Which error categories are easier to predict? Where do non-reasoning models perform notably worse? We examine this in §5.1.6.
@@ 5.1.1 Task Difficulty - Context Length and Generation Horizon
As seen in Table 2, the distribution of raw JSON input token lengths injested to perform our task cuts close to the input context limit of several LLMs - with the maximum input trace length always being twice longer than the input length limit, and even the mean itself sometimes going over. Furthermore, even the typical output token length horizon the LLMs need to generate for the task exceeds the 1K tokens on average, with the maximum being $\approx$3.7K at the least. Besides being a significant % of the maximum output length, this indicates the difficultly long generation horizon TRAIL needs.
@@
Table 3: Correlations b/w Input Length & Performance
5.1.2 Long Context Ability and Model Performance
We compare how the models in Table 1 rank based on their aggregate performance on TRAIL vis-a-vis the relative ranking of the subsets of these models that occur on updated long-context benchmark leaderboards Longbenchv2 and fiction.live’s LongContextBench Bai et al. (2024b); Ficlive (2025), and notice this differs for only one model (o3 being third best rather than best on the latter). We defer the complete detail of these rankings to §A.2.
5.1.3 Performance vs Input Length
We find all performance metrics to be anti-correlated with input length, as detailed in Table 3. This supports the hypothesis that longer input raw traces increase the difficulty of TRAIL for models.
@@ 5.1.4 Reasoning vs Non-Reasoning Models
From Table 1, we see all reasoning models except o1 outperforming non-reasoning ones on both Error category F1 and Location Accuracy. On Joint Accuracy, the gap between the two families is larger — Reasoning models other than o1 perform at 1.5-8 times the best performing non-reasoning model.
@@ 5.1.5 Does Reasoning Effort Matter?
To systematically assess the impact of reasoning extent, we experiment with the same model (o3) at "high," "medium," and "low" reasoning effort levels, as set by OpenAI’s reasoning.effort parameter. We find that all three metrics, including Category F1 (0.296 → 0.277 → 0.264), decrease as reasoning effort decreases. These results empirically support that TRAIL performance benefits from higher reasoning effort at test time, and that the superior results for reasoning models are not solely due to improved pre- or post-training (§5.1.4). Full ablation results are in Appendix §A.4.
@@ 5.1.6 Performance Across Categories
Figure 4: Heatmap for Error Category F1 across models; categories are ordered left to right based on their support
@@
Among the most challenging categories, Context Handling Failures stand out, as nearly all models score an F1 of 0.00, indicating these errors demand advanced reasoning. The only exception is claude-3.7-sonnet, which achieves a relatively better score of 0.18. Tool Selection Errors are also difficult to predict, with most models scoring between 0.00 and 0.08, apart from gemini-2.5-pro (0.26), claude-3.7-sonnet (0.27), and especially o3 (0.53), suggesting this is a complex error type. Similarly, Task Orchestration shows uniformly low scores across models (0.00–0.08) except for gemini-2.5-flash, which stands out with a much higher F1 of 0.47.
@@
There are also categories where model performance diverges interestingly. For Goal Deviation, gemini-2.5-pro and gemini-2.5-flash perform best (0.70 and 0.41, respectively), while claude-3.7-sonnet and o3 perform moderately (0.31, 0.24); o1 and other non-reasoning models score the lowest ($\leq$ 0.05). In the case of Poor Information Retrieval, the two Gemini models are again notably better (0.50 and 0.53), with others at <0.30, suggesting better diagnosis of failures related to context.
@@
Language-Only errors, a subtype of hallucination, are detected relatively well by all models (0.14–0.59), implying that these are easier for models to predict even without advanced reasoning capabilities. For Formatting Errors, performance is non-monotonic: gpt-4.1 (0.43) and the gemini-2.5 models (0.44–0.57) perform well, while o1, o3, and claude-3.7-sonnet perform worse (0.23–0.31). It is notable that o1 and gpt-4.1 outscore o3 on this category, despite being older and non-reasoning respectively. We defer some model-specific observations to §A.6
@@ 6 Conclusion
In this work, TRAIL, a new taxonomy for classifying agentic errors, along with an expert-curated dataset of 148 agentic problem instances and 841 unique errors from GAIA and SWE Bench. Current SOTA models perform poorly as LLM Judges on this dataset, with Gemini 2.5-pro achieving only 18% joint accuracy on GAIA and 5% on SWE Bench; three out of eight models cannot even process the full context. These results highlight that existing models struggle to systematically evaluate complex agentic traces, due to the inherent complexity of agentic systems and LLM context limitations. A new framework is needed for scalable, systematic evaluation of agentic workflows.
Limitations
The TRAIL dataset and taxonomy are primarily focused on text-only inputs and outputs but recent advancements in multimodal agentic systems require careful extension of the taxonomy to handle errors arising from new categories such as multimodal tool use. One additional limitation of TRAIL is the large number of tail categories with very few examples. It is important to ensure correctness of LLM-Judges on these categories due to the high-impact nature of the failures. Future research work can look into synthetic data generation for high-impact, low-occurrence categories by systematically modifying existing traces to induce catastrophic irrecoverable failures within the LLM context.
@@
While curating this dataset, we ensure that annotators are only selected based on their age (18+) and their expertise in the computer science field. Annotator selection was not based on nationality, language, gender or any other characteristic apart from these two criteria. We pay annotators a total of \$12.66 per trace where each trace takes 30-40 minutes to annotate. We ensure that the traces do not contain any PII or any explicit or biased content by manually verifying traces before forwarding these to annotators. The annotators were made aware of the open-sourcing of their work and consent was obtained beforehand.
@@ A.1 Prompt Structure
A.2 Long Context Leaderboard Rankings vs TRAIL
@@
A.3 Evaluation Setup
A.4 Reasoning Effort Ablations
@@ A.5 Span Statistics
A.6 Model-Specific Observations
gemini-2.5-pro is clearly the strongest overall, excelling particularly at Goal Deviation (0.70), Poor Information Retrieval (0.50), Tool Output Misinterpretation (0.67), and Environment Setup Errors (0.57). By contrast, gpt-4.1 shows great variability, performing very well or moderately on some categories such as Instruction Non-compliance, Language-only, Formatting Errors, and Resource Abuse, but dipping below 0.10 or even hitting zero on others, including Goal Deviation, Tool Selection Errors, Task Orchestration, Tool-related Hallucinations, and Context Handling Failures.
A.6.1 Visualizing Token Length Distributions
(a) GAIA
(b) SWEBench
A.7 Complete Data Annotation, Validation and Agreement Details
Due to the large trace size—often exceeding LLM context limits (§5.1.1)—we conducted four independent verification rounds with ML researchers for quality assurance. Annotators evaluated each LLM and tool span in sequence, marking span ID, error category, evidence, description, and impact (Low/Medium/High) per our taxonomy, and rated overall traces for instruction adherence, plan optimality, security, and reliability (see § A.7.1). On average, annotating a GAIA trace took 30 minutes, SWE Bench traces 40 minutes, and verification added  20 minutes, totaling $\approx$110 and 120 minutes per trace, respectively [^5] We did not explore and verify information (web-based or otherwise) from contents external to the trace because our baseline models are not expected to do so. Verifying such information will add more time to this estimate.. For SWE Bench, 30 traces (444 spans) were reviewed, with 5.63% of spans modified—mainly Resource Abuse (33.33%), Language-only Hallucinations (20.83%), and Tool-related Hallucinations (12.5%). For GAIA, 33 traces (697 spans) were reviewed, with 5.31% revised, primarily Language-only Hallucinations (23.08%), Resource Abuse (19.23%), and Poor Information Retrieval (19.23%). These results indicate high inter-annotator agreement during curation.
A.7.1 Annotation Rubric
@@
A.8 Correlation scores for Rubrics
@@
A.9 Distribution of Impact Levels in TRAIL instances
@@
A.10 Agent Orchestrations for TRAIL